# ReflectDiffu: Reflect between Emotion-intent Contagion and Mimicry for Empathetic Response Generation via a RL-Diffusion Framework

### v1

@@ Abstract
Jiahao Yuan1, Zixiang Di1, Zhiqing Cui2, Guisong Yang3*, Usman Naseem4* 1East China Normal University 2Nanjing University of Information Science and Technology 3University of Shanghai for Science and Technology 4Macquarie University {51275900024,51265901113}@stu.ecnu.edu.cn, 202383090003@nuist.wdu.cn, gsyang@usst.edu.cn, usman.naseem@mq.eu.au
@@ 1 Introduction
  • We utilize LLMs to expand annotation on EmpatheticDialogue Rashkin et al. (2018) and create a multi-task dataset for emotional reason masking, emotion prediction, intent prediction, and empathetic dialogue generation. Our dataset is available [^1] https://anonymous.4open.science/status/ReflectDiffu-4C12.
@@ 3.3 Multi-task Rational Response Generation Decoder
$R(e)=\begin{cases}\texttt{sigmoid}(\textit{Emo}_{\textit{pos}}[i]\cdot\textit{intent}_{\textit{refer}})&\textit{if }\textit{is\_pos}(e)\\ \texttt{sigmoid}(\text{Emo}_{\textit{neg}}[i]\cdot\textit{intent}_{\textit{refer}})&\textit{otherwise}\end{cases}.$ (9)
@@
$\textit{L}_{\textit{twice}}=\text{L}_{\text{kl}_{\text{pos}}}+\text{L}_{\text{kl}_{\text{neg}}}+\text{L}_{\text{intent}}.$ (10)
@@ 3.4 Training
Lastly, all parameters of ReflectDiffu are trained jointly in an end-to-end manner to optimize the model by integrating all losses L, using loss weight averaging as follows:
@@ B.2 Case Study in Misclassification
We deliberately selected a ReflectDiffu emotion recognition error case to validate the effectiveness of our reflection mechanism. Table 5 compares responses from various models, including MOEL, MIME, EmpDG, KEMP, CASE, CAB, ChatGLM-6B, and ReflectDiffu, to a user’s context of feeling hopeful after applying for graduate school. Initially, ReflectDiffu misidentified the emotion as "joyful" and the intent as "acknowledging." However, after employing the reflection mechanism, it correctly identified the intent as "encouraging." This demonstrates the model’s capability to correct errors and generate more empathetic responses through its reflection mechanism.
To further illustrate the effectiveness of the reflection mechanism, we visualized the changes in emotion-intent selection dynamics using a series of plots.

### v2

@@ Abstract
Jiahao Yuan1, Zixiang Di1, Zhiqing Cui2, Guisong Yang3*, Usman Naseem4* 1East China Normal University 2Nanjing University of Information Science and Technology 3University of Shanghai for Science and Technology 4Macquarie University {51275900024,51265901113}@stu.ecnu.edu.cn, 202383090003@nuist.edu.cn, gsyang@usst.edu.cn, usman.naseem@mq.edu.au
@@ 1 Introduction
  • We utilize LLMs to expand annotation on EmpatheticDialogue Rashkin et al. (2018) and create a multi-task dataset for emotional reason masking, emotion prediction, intent prediction, and empathetic dialogue generation. Our dataset is available [^1] https://anonymous.4open.science/r/ReflectDiffu-4C12.
@@ 3.3 Multi-task Rational Response Generation Decoder
$R(e)=\begin{cases}\texttt{sigmoid}(\textit{Emo}_{\textit{pos}}[i]\cdot\textit{intent}_{\textit{refer}})&\textit{if }\textit{is\_pos}(e)\\ \texttt{sigmoid}(\textit{Emo}_{\textit{neg}}[i]\cdot\textit{intent}_{\textit{refer}})&\textit{otherwise}\end{cases}.$ (9)
@@
$\textit{L}_{\textit{twice}}=\textit{L}_{\textit{kl}_{\textit{pos}}}+\textit{L}_{\textit{kl}_{\textit{neg}}}+\textit{L}_{\textit{intent}}.$ (10)
@@ 3.4 Training
Lastly, all parameters of ReflectDiffu are trained jointly in an end-to-end manner to optimize the model by integrating all losses L, employing loss weight averaging with hyperparameters $\delta,\zeta,\eta$ as follows:
@@ B.1 Explanation of the hyperparameter $n$ of $\text{intent}_{\text{infer}}$
Qwen2-7B+COT That’s wonderful! Applying for graduate school is a significant step
llama3.1-8B+COT Congratulations on taking this important step! That’s fantastic!
@@ B.2 Case Study in Misclassification
We deliberately selected a ReflectDiffu emotion recognition error case to validate the effectiveness of our reflection mechanism. Table 5 compares responses from various models, including MOEL, MIME, EmpDG, KEMP, CASE, CAB, Qwen2-7B+COT,llama3.1-8B+COT and ReflectDiffu, to a user’s context of feeling hopeful after applying for graduate school. Initially, ReflectDiffu misidentified the emotion as "joyful" and the intent as "acknowledging." However, after employing the reflection mechanism, it correctly identified the intent as "encouraging." This demonstrates the model’s capability to correct errors and generate more empathetic responses through its reflection mechanism.